OpenStudy (anonymous):

Jeremy is building a large deck for a community center. The deck is shaped as a rectangle. The width of the deck is 29 feet. The perimeter of the deck is to be at least 134 feet. a. Write an inequality that represents all possible values for the length of the deck. b. Find all possible values for the length of the deck

OpenStudy (radar):

The width is fixed at 29 feet. It is to be a rectangle, so both widths will be 29 feet. That means 58 feet is reserved for the width. Now the perimeter is to be at least 134 feet. To me that means the perimeter is at a minimum 134 ft, but can be larger. Is that how you understand it?

OpenStudy (radar):

Perimeter P=2w + 2L P = 58 + 2L\[P \ge 134\] \[2L + 58 \ge 134\]That would be part A, but it can be further simplified.

OpenStudy (radar):

Express the inequality in terms of L (length) not 2L.

OpenStudy (radar):

No monologue for me.